# Provably Efficient Reinforcement Learning with Multinomial Logit Function Approximation

### v1

@@ Appendix I Proof of Theorem 4
Our proof is similar to the proof of the lower bound for linear mixture MDPs in Zhao et al. [2023]. We prove this lemma by translating this problem to learning $H/2$ logistic bandit problems.
@@
$\mathbb{P}_{i,1}(s_{i,2}^{*}\mid s_{i,1},a)$ $=\frac{\exp(\phi(s_{i,2}^{*}\>|\>s_{i,1},a)^{\top}\theta_{i,1}^{*})}{1+\exp(\phi(s_{i,2}^{*}\>|\>s_{i,1},a)^{\top}\theta_{i,1}^{*})}=\rho_{a},\quad\mathbb{P}_{i,1}(s_{i,2}\mid s_{i,1},a)=1-\rho_{a}.$
For the second layer, $\forall s=s_{i,2}^{*},s_{i,2}$, and $a\in\mathcal{A}$, $\mathbb{P}_{i,2}(s_{i,3}\mid s,a)=1$. The reward satisfies $r_{k}(s_{i,1},a)=0$ for the first layer and $r_{k}(s_{i,2}^{*},a)=1$, $r_{k}(s_{i,2},a)=0$ for the second layer for all $a\in\mathcal{A}$.
Consider the logistic bandit problem where a learner selects action $x\in\mathbb{R}^{d}$ and receives a reward $r_{k}$ sampled from Bernoulli distributed with mean $\mu(x^{\top}\theta^{*})=(1+\exp(-x^{\top}\theta^{*})^{-1})$. Thus, learning in this block of MDP can be regarded as learning a $d$-dimensional logistic bandit problem with $A$ arms, where the arm set is $\phi(s_{i,2}^{*}\>|\>s_{i,1},a)$ and the expected reward of each arm is $\rho_{a}$. It is also clear that the optimal policy at state $s_{i,1}$ is to select action $a_{i,1}^{*}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\phi(s_{i,2}^{*}\>|\>s_{i,1},a)$. Thus, learning this MDP is equal to learning $H/2$ logistic bandit problems. This finishes the proof. ∎
Appendix J Supporting Lemmas
In this part, we provide several supporting lemmas used in the proofs of the main results.
@@
Let $\left\{\mathcal{F}_{t}\right\}_{t=0}^{\infty}$ be a filtration. Let $\left\{\eta_{t}\right\}_{t=1}^{\infty}$ be a real-valued stochastic process such that $\eta_{t}$ is $\mathcal{F}_{t}$-measurable and $\eta_{t}$ is conditionally zero-mean $R$-sub-Gaussian for some $R\geq 0$ i.e. $\forall\lambda\in\mathbb{R}$, $\mathbb{E}\left[e^{\lambda\eta_{t}}\mid F_{t-1}\right]\leq\exp\left({\lambda^{2}R^{2}}/{2}\right)$. Let $\left\{X_{t}\right\}_{t=1}^{\infty}$ be an $\mathbb{R}^{d}$-valued stochastic process such that $X_{t}$ is $\mathcal{F}_{t-1}$-measurable. Assume that $V\in\mathbb{R}^{d\times d}$ is a positive definite matrix. For any $t\geq 0$, define $V_{t}=V+\sum_{s=1}^{t-1}X_{s}X_{s}^{\top},S_{t}=\sum_{s=1}^{t-1}\eta_{s}X_{s}$. Then, for any $\delta>0$, with probability at least $1-\delta$, for all $t\geq 0$,
$\left\|S_{t}\right\|_{V_{t}^{-1}}^{2}\leq 2R^{2}\log\left(\frac{\det\left(V_{t}\right)^{1/2}\det(V)^{-1/2}}{\delta}\right).$
@@
Suppose $x_{1},\ldots,x_{t}\in\mathbb{R}^{d}$ and for any $1\leq s\leq t$, $\lVert x_{s}\rVert_{2}\leq L$. Let $V_{t}=\lambda I_{d}+\sum_{s=1}^{t}x_{s}x_{s}^{\top}$ for $\lambda\geq 0$. Then, we have $\det(V_{t})\leq(\lambda+{tL^{2}}/{d})^{d}$.
@@
Let $Z$ be a positive define matrix and $\mathcal{X}$ be a convex set, define $\widetilde{\mathbf{x}}=\operatorname*{arg\,min}_{\mathbf{x}\in\mathcal{X}}\langle\mathbf{g},\mathbf{x}\rangle+\frac{1}{2\eta}\lVert\mathbf{x}-\widehat{\mathbf{x}}\rVert_{Z}^{2}$. Then, for all $\mathbf{x}^{*}\in\mathcal{X}$, we have
$\lVert\widetilde{\mathbf{x}}-\widehat{\mathbf{x}}\rVert_{Z}\leq 2\eta\lVert\mathbf{g}\rVert_{Z^{-1}}\quad\mbox{and}\quad\langle\mathbf{g},\widehat{\mathbf{x}}-\mathbf{x}^{*}\rangle\leq\frac{\eta}{2}\lVert\mathbf{g}\rVert_{Z^{-1}}^{2}+\frac{1}{2\eta}\lVert\widehat{\mathbf{x}}-\mathbf{x}^{*}\rVert_{Z}^{2}-\frac{1}{2\eta}\lVert\widetilde{\mathbf{x}}-\mathbf{x}^{*}\rVert_{Z}^{2}.$
@@
Define $Q:\mathbb{R}^{K}\rightarrow\mathbb{R}$, such that for any $\mathbf{u}=\left(u_{1},\ldots,u_{K}\right)\in\mathbb{R}^{K},Q(\mathbf{u})=$ $\sum_{i=1}^{K}\frac{\exp\left(u_{i}\right)}{v+\sum_{k=1}^{K}\exp\left(u_{k}\right)}$. Let $p_{i}(\mathbf{u})=\frac{\exp\left(u_{i}\right)}{v+\sum_{k=1}^{K}\exp\left(u_{k}\right)}$. Then, for all $i\in[K]$,

### v2

@@ Appendix H Proof of Theorem 3
Our proof is similar to the proof of the lower bound for adversarial linear mixture MDPs with the unknown transition in Zhao et al. [2023]. We prove this lemma by reducing the MNL mixture MDP problem to a sequence of binary logistic bandit problems.
@@
$\mathbb{P}_{i,1}(s_{i,2}^{*}\mid s_{i,1},a)$ $=\frac{\exp(\phi(s_{i,2}^{*}\>|\>s_{i,1},a)^{\top}\theta_{i,1}^{*})}{1+\exp(\phi(s_{i,2}^{*}\>|\>s_{i,1},a))^{\top}\theta_{i,1}^{*})}=\rho_{a},$
$\quad\mathbb{P}_{i,1}(s_{i,2}\mid s_{i,1},a)$ $=\frac{1}{1+\exp(\phi(s_{i,2}^{*}\>|\>s_{i,1},a)^{\top}\theta_{i,1}^{*})}=1-\rho_{a}.$
For the second layer, the MDP instance satisfies that $\forall s=s_{i,2}^{*},s_{i,2}$, and $a\in\mathcal{A}$, $\mathbb{P}_{i,2}(s_{i,3}\mid s,a)=1$. The reward function satisfies $r_{k}(s_{i,1},a)=0$ for the first layer and $r_{k}(s_{i,2}^{*},a)=1$, $r_{k}(s_{i,2},a)=0$ for the second layer for all $a\in\mathcal{A}$.
Consider the logistic bandit problem where a learner selects action $x\in\mathbb{R}^{d}$ and receives a reward $r_{k}$ sampled from Bernoulli distributed with mean
$\mu(x^{\top}\theta^{*})=\frac{1}{1+\exp(x^{\top}\theta^{*})}.$
By this configuration, we can see that learning in this block of MDP can be regarded as learning a $d$-dimensional logistic bandit problem with $A$ arms, where the arm set is $\phi(s_{i,2}^{*}\>|\>s_{i,1},a)$ and the expected reward of each arm is $\rho_{a}$. It is also clear that the optimal policy at state $s_{i,1}$ is to select action $a_{i,1}^{*}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\phi(s_{i,2}^{*}\>|\>s_{i,1},a)$. Thus, learning this MNL mixture MDP equals to learning $H/2$ logistic bandit problems. This finishes the proof. ∎
Appendix I Supporting Lemmas
In this section, we provide several supporting lemmas used in the proofs of the main results.
@@ Appendix I Supporting Lemmas
Let $\left\{\mathcal{F}_{t}\right\}_{t=0}^{\infty}$ be a filtration. Let $\left\{\eta_{t}\right\}_{t=1}^{\infty}$ be a real-valued stochastic process such that $\eta_{t}$ is $\mathcal{F}_{t}$-measurable and $\eta_{t}$ is conditionally zero-mean $R$-sub-Gaussian for some $R\geq 0$ i.e. $\forall\lambda\in\mathbb{R}$, $\mathbb{E}\left[e^{\lambda\eta_{t}}\mid F_{t-1}\right]\leq\exp\left({\lambda^{2}R^{2}}/{2}\right)$. Let $\left\{X_{t}\right\}_{t=1}^{\infty}$ be an $\mathbb{R}^{d}$-valued stochastic process such that $X_{t}$ is $\mathcal{F}_{t-1}$-measurable. Assume that $V$ is a $d\times d$ positive definite matrix. For any $t\geq 0$, define
$V_{t}=V+\sum_{s=1}^{t-1}X_{s}X_{s}^{\top}\quad S_{t}=\sum_{s=1}^{t-1}\eta_{s}X_{s}.$
Then, for any $\delta>0$, with probability at least $1-\delta$, for all $t\geq 0$,
$\left\|S_{t}\right\|_{V_{t}^{-1}}^{2}\leq 2R^{2}\log\left(\frac{\operatorname{det}\left(V_{t}\right)^{1/2}\operatorname{det}(V)^{-1/2}}{\delta}\right).$
@@
Suppose $x_{1},\ldots,x_{t}\in\mathbb{R}^{d}$ and for any $1\leq s\leq t$, $\lVert x_{s}\rVert_{2}\leq L$. Let $V_{t}=\lambda I_{d}+\sum_{s=1}^{t}x_{s}x_{s}^{\top}$ for $\lambda\geq 0$. Then, for any $1\leq s\leq t$, we have
$\det(V_{t})\leq\left(\lambda+\frac{tL^{2}}{d}\right)^{d}.$
@@
Let $Z$ be a positive define matrix and $\mathcal{X}$ be a convex set,
$\widetilde{\mathbf{x}}=\operatorname*{arg\,min}_{\mathbf{x}\in\mathcal{X}}\langle\mathbf{g},\mathbf{x}\rangle+\frac{1}{2\eta}\lVert\mathbf{x}-\widehat{\mathbf{x}}\rVert_{Z}^{2}.$
Then, for all $\mathbf{x}^{*}\in\mathcal{X}$, we have
$\lVert\widetilde{\mathbf{x}}-\widehat{\mathbf{x}}\rVert_{Z}\leq 2\eta\lVert\mathbf{g}\rVert_{Z^{-1}}.$
and
$\langle\mathbf{g},\widehat{\mathbf{x}}-\mathbf{x}^{*}\rangle\leq\frac{\eta}{2}\lVert\mathbf{g}\rVert_{Z^{-1}}^{2}+\frac{1}{2\eta}\lVert\widehat{\mathbf{x}}-\mathbf{x}^{*}\rVert_{Z}^{2}-\frac{1}{2\eta}\lVert\widetilde{\mathbf{x}}-\mathbf{x}^{*}\rVert_{Z}^{2}.$
@@
Define $Q:\mathbb{R}^{K}\rightarrow\mathbb{R}$, such that for any $\mathbf{u}=\left(u_{1},\ldots,u_{K}\right)\in\mathbb{R}^{K},Q(\mathbf{u})=$ $\sum_{i=1}^{K}\frac{\exp\left(u_{i}\right)}{v+\sum_{k=1}^{K}\exp\left(u_{k}\right)}$. Let $p_{i}(\mathbf{u})=\frac{\exp\left(u_{i}\right)}{v+\sum_{k=1}^{K}\exp\left(u_{k}\right)}$. Then, for all $i\in[K]$, we have